# IPO: Iterative Preference Optimization for Text-to-Video Generation

### v3

@@ Abstract
Video foundation models have achieved significant advancement with the help of network upgrade as well as model scale-up. However, they are still hard to meet requirements of applications due to unsatisfied generation quality. To solve this problem, we propose to align video foundation models with human preferences from the perspective of post-training in this paper. Consequently, we introduce an Iterative Preference Optimization strategy to enhance generated video quality by incorporating human feedback. Specifically, IPO exploits a critic model to justify video generations for pairwise ranking as in Direct Preference Optimization or point-wise scoring as in Kahneman-Tversky Optimization. Given this, IPO optimizes video foundation models with guidance of signals from preference feedback, which helps improve generated video quality in subject consistency, motion smoothness and aesthetic quality, etc. In addition, IPO incorporates the critic model with the multi-modality large language model, which enables it to automatically assign preference labels without need of retraining or relabeling. In this way, IPO can efficiently perform multi-round preference optimization in an iterative manner, without the need of tediously manual labeling. Comprehensive experiments demonstrate that the proposed IPO can effectively improve the video generation quality of a pretrained model and help a model with only 2B parameters surpass the one with 5B parameters. Besides, IPO achieves new state-of-the-art performance on VBench benchmark.The source code and model checkpoints, along with implementation details and resources, are openly available at this GitHub repository.
Figure 1: Comparison of (a) traditional single-round preference optimization and (b) the proposed Iterative Preference Optimization for aligning video foundation model with human feedback. IPO introduces a critic model to automatically labeling video data without need of tedious manual efforts, making iterative update possible in an efficient way.
Figure 2: The overview of our proposed Iterative Preference Optimization framework. IPO consists of three major parts: (a) Human Preference Dataset. It is used to train a critic model. This dataset only requires one-round manual annotation without need of relabeling in the iterative framework; (b) Critic Model. It is learned to automatically annotate generated videos with preference labels, which help eliminate tedious manual efforts for multi-round optimization; (c) Iterative Optimization. It incorporates the critical model to iteratively optimize video foundation model. In this way, IPO can efficiently tune the base model to align with human preference and enhance generation ability in aspects of subject consistency, motion smoothness and aesthetic quality, etc. Best viewed in 2x zoom.
@@ 1 Introduction
Text-to-video generation has drawn lots of attention due to its great potentials in movie products, video effects and user-generated contents, etc. Recent works [2, 16, 10] introduce the Transformer architecture into diffusion models and improve the video generation quality by simply scaling up the parameter size, e.g., CogVideoX-5B [43], Mochi-10B [35] and HunyuanVideo-13B [15]. Although achieving impressive performance, they still face challenges for generating user-satisfied videos with consistent subject, smooth motion and high-aesthetic picture. In addition, their large model sizes make them slow to produce videos. These drawbacks limit their applications in practice.
To align generation results with human preference, post-training techniques have shown their effectiveness in fields of Large Language Models (LLMs) and Text-to-Image models (T2Is). For instance, Direct Preference Optimization (DPO) [27] exploits pairwisely ranked data to make LLMs know the language style liked by human. While Kahneman-Tversky Optimization (KTO) [7] uses pointwisely binary data to reinforce T2Is to learn how to maximumly align the objective with human expectation. These techniques promote generative models to derive user-satisfied contents and give us inspiration for improving video generation models.
Thus, we propose to align Text-to-Video models (T2Vs) with human preference from the perspective of post-training in this paper. In particular, we present the Iterative Preference Optimization (IPO) framework for enhancing T2Vs. IPO introduces preference priors for finetuning T2Vs in the form of reinforcement learning, which considers subject consistency, motion smoothness and aesthetic quality, etc. This leads to improved video generations that are well aligned with human preference. In addition, as show in Fig. 1, difference from prior single-round methods, IPO adopts a multi-round optimization strategy to reinforce T2Vs in an iterative manner, which strengths the models continuously. In this way, IPO can effectively improve T2Vs without need of large-scale dataset and expensive supervised finetuning.
Specifically, IPO consists of three core parts as shown Fig. 2: the preference dataset, the critic model and the iterative learning framework. IPO first collects a preference dataset for training the critic model. To get this dataset, IPO predefines several categories, including human, animal, action, etc. Then, IPO randomly combines these categories and uses LLM to extend them as prompts for generating videos with T2V models. Given these videos, IPO manually annotates them with two kinds of labels: one is pairwise ranking label that depicts preference comparison of two videos; another is pointwise scoring label that describes video quality as good or bad in a binary manner. After collecting the preference dataset, IPO uses it to train a critic model for justifying the quality of input videos. Concretely, IPO exploits Multi-modality Large Language Models (MLLMs) to achieve the above goal considering their powerful capability for recognizing video contents. In addition, IPO combines instruction-tuning to learn the critic model, making it able to generate pairwise and pointwise labels with different prompts. Given the critic model, IPO plays as a plugin for preference optimization of T2Vs without requirements of retraining or relabeling. For a specific T2V model, IPO can either use Diffusion-DPO [36] with pairwise ranking data or Diffusion-KTO [18] with pointwise scoring data. In addition, IPO adopts the critic model for automatically labeling videos from the reinforced T2Vs, which can be used as prior information for further improving T2Vs. In the iterative way, IPO can consistently improve T2Vs for generating high-quality videos.
Comprehensive experiments on benchmark VBench demonstrate that the proposed IPO can effectively and efficiently reinforce T2Vs to produce user-satisfied videos with consistent subject, smooth motion as well as high-aesthetic quality. In addition, given CogVideoX as the video foundation model, IPO can improve a 2B model to surpass the 5B counterpart. Our contributions can be summarise into three folds: First, we introduce a new preference optimization framework into text-to-video generation, which well align generations with human preference. Second, we present a novel iterative post-training framework for enhancing foundation models in a multi-round manner, differing from prior single-round methods. Third, we set new state-of-the-art on the benchmark VBench both in quantity and quality. To advance future researches in the area of video generation, we will release the preference dataset, the codes, the critic models as well as T2V models.
@@ 2 Related Work
Text-to-video generation Recent advances in video generation have been largely driven by diffusion models [11, 31, 32, 46], achieving remarkable improvements in diversity, fidelity, and quality [4, 6, 13, 20, 29, 38, 40]. Scaling model size and training data further enhance performance [12, 30, 39]. In Text-to-Video (T2V) generation, pre-training from scratch requires substantial computational resources and large-scale datasets [8, 22, 16, 2, 15]. Current methods [1, 37, 9] extend text-to-image models by incorporating spatial and temporal attention modules, often using joint image-video training [23], with [9] proposing a plug-and-play temporal module for personalized image models. State-of-the-art pixel quality and temporal consistency have been achieved by [47, 3], while diffusion models based on the DiT structure [2, 43, 3] enhance spatiotemporal coherence and scalability, highlighting the potential of Transformer-based architectures. Despite of their success, they still cannot meet requirements in real scenarios due to lack of alignment with human preference.
Post-training for human preference alignment Post-training is a common technique exploited in Large Language Models (LLMs)s [24, 28, 50]. Reinforcement Learning from Human Feedback (RLHF) [24] learns a reward function from comparison data to align the policy model. Rank Responses for Human Feedback (RRHF) [45] extends supervised fine-tuning, while Sequence Likelihood Calibration (SLiC) [49] leverages offline RL data. Direct Preference Optimization (DPO) [27] improves stability and efficiency by directly optimizing human preference data, and Kahneman-Tversky Optimization (KTO) [7] aligns models via a prospect-theoretic approach. Rejection Sampling Optimization (RSO) [21] enhances alignment using statistical rejection sampling. Prior works have introduced this technique into enhancement of image foundation models, such as [5, 25, 41]. Diffusion-DPO [36] extends DPO for diffusion models, optimizing model likelihood, while Diffusion-KTO [18] maximizes expected human utility without costly pairwise preference data. The usage of post-training in video generation is still rare. Current works mainly focus on the design of reward models. [44] introduced a temporal decaying reward (TAR) mechanism, prioritizing central frames for stable and coherent video generation. Meanwhile, [17] optimized ordinary differential equation (ODE) solvers by exploring energy function designs, enhancing T2V training through motion prior extraction. VADER [26] fine-tunes video diffusion models using reward model gradients and pre-trained vision models, achieving strong generalization even for unseen cues, thereby improving T2V adaptability and performance. However, the above strategies can only achieve minor improvements and not fully explore the potential of post-training.
Figure 3: The distribution statistics of Human Preference Dataset on prompts in (a) and (b) as well as scoring/ranking in (c) and (d).
@@ 3 METHOD
The proposed Iterative Preference Optimization framework consists of three major parts: (1) A human preference dataset for training critic model; (2) A critic model for automatically labeling generated videos; (3) An iterative optimization process for enhancing the base text-to-video models. Fig. 2 illustrates the whole pipeline of IPO. We will explain details of each part in the following sections.
3.1 Human Preference Dataset
Video-text Collection
The generated video dataset is created using the CogVideoX-2B model [43]. We generate diverse prompts covering various character types, which are then combined and refined using the Qwen2.5 [42]. To increase diversity, we adjust parameters like temperature, top-k, and top-p. As shown in Fig. 3, the dataset includes over ten categories, such as Human, Architecture, Vehicles, and Animals, as well as more than ten different actions of human and animal. For each prompt, the video generation model produces four variants with different random seeds. The dataset is iteratively expanded through an optimized regeneration strategy to improve preference alignment.
Human Annotation When aligning human feedback, we focus on three dimensions, inspired by the text-to-image evaluation work Evalalign [34]: consistency between text and video, video faithfulness, and additional motion smoothness. To capture human preferences, we employ two annotation methods: Pairwise Ranking and Pointwise Scoring. For pairwise ranking, annotators are shown two videos generated from the same prompt and asked to evaluate them based on predefined criteria. They are required to: (1) provide detailed natural language feedback, and (2) identify specific quality issues in each video. The annotated comparisons are then organized to create training samples.
{<Video1>, <Video2>, Question,Evaluation Dimension,Ranking $\left\{\begin{array}[]{l}\texttt{<Video1>, <Video2>, Question,}\\ \texttt{Evaluation Dimension,}\\ {\color[rgb]{0,0,1}\text{Ranking~{}Answer}},{\color[rgb]{0,0,1}\text{Reason}}\end{array}\right\}$
For pointwise scoring, each video is assessed across three dimensions, with scores categorized into three levels: Good, Normal, and Poor. Similar to the ranking process, annotators provide detailed justifications for their scores to ensure transparency and consistency. The annotated evaluations are then organized to create training samples.
{<Video>, Question,Evaluation Dimension,Scoring $\left\{\begin{array}[]{l}\texttt{<Video>, Question,}\\ \texttt{Evaluation Dimension,}\\ {\color[rgb]{0,0,1}\text{Scoring~{}Answer}},{\color[rgb]{0,0,1}\text{Reason}}\end{array}\right\}$
Each dataset is annotated by at least three individuals. The final annotation is determined through voting, selecting the one with the highest consistency among reviewers.
3.2 Critic Model Training
Large Visual-Language Models (VLMs) exhibit strong generalization across diverse tasks due to pre-training on extensive multimodal data. Their capability to capture complex text-image relationships makes them well-suited for critic modeling, where alignment with human preferences is crucial. Leveraging VLMs as critic models enhances video generation by improving preference predictions through their robust generalization ability.
Fig. 2 illustrates our supervised fine-tuning process. The training starts by passing video frames through an image encoder to extract image tokens. These tokens, together with the textual caption, question, and answer, are then fed into a pre-trained Visual-Language Model (VLM). Formally, each training sample consists of a triplet: multimodal input $M$, which includes both the video and its corresponding textual description; question $Q$; and answer $A$.
The optimization objective is derived from the autoregressive loss function commonly used in LLM training, but applied specifically to the answer as:
$L(\phi)=-\sum^{N}_{i=1}\log p(A_{i}\mid Q,M,A_{<i};\phi),$ (1)
where $N$ is the length of the ground-truth answer. This loss encourages the model to get answers that align with provided questions and multimodal context. Depending on the method, the model receives different inputs. In the pointwise case, it evaluates a single video-text pair. In the pairwise case, it compares two pairs to determine which aligns better with the text. The optimization focuses on individual quality in pointwise and relative comparison in pairwise.
3.3 T2V Alignment
3.3.1 Alignment Optimization
In Reinforcement Learning with Human Feedback (RLHF), the primary objective is to optimize the model to maximize the expected reward given context $c$. This is achieved by incorporating a regularization term that penalizes deviations from a reference distribution, ensuring the optimized model remains consistent with prior knowledge or previously trained models. The optimization objective is as:
@@ 3.3.1 Alignment Optimization
$\max_{\pi_{\theta}}\mathbb{E}_{x_{0}\sim\pi_{\theta}(x_{0}|c)}\left[r(x_{0},c)-\beta\log\frac{\pi_{\theta}(x_{0}|c)}{\pi_{ref}(x_{0}|c)}\right],$ (2)
where $r(x_{0},c)$ denotes the reward associated with a state $x_{0}$ given context $c$, $\pi_{\theta}(x_{0}|c)$ is the policy of the model parameterized by $\theta$, and $\pi_{ref}(x_{0}|c)$ is the reference policy. The hyperparameter $\beta$ controls the strength of the regularization, typically set as a fixed constant. This objective function aims to balance reward maximization with regularization, which prevents the model from straying too far from the reference distribution, ensuring stability and generalization [48]. And the loss function for the policy $\pi_{\theta}(\cdot|c)$ is formulated in terms of the expected value over the reward and the KL-divergence penalty as following:
$\mathbb{E}_{x_{0}\sim\pi_{\theta}(x_{0}|c)}\Big{[}-r(x_{0},c)-\beta\log\frac{\pi_{ref}(x_{0}|c)}{\pi_{\theta}(x_{0}|c)}\Big{]}=\\ \beta\mathbb{KL}\left(\pi_{\theta}(x_{0}|c)\Big{\|}\pi_{ref}(x_{0}|c)\exp\left(\frac{1}{\beta}r(x_{0},c)\right)\right),$ (3)
Minimizing this loss leads to optimal policy $\pi^{*}(x_{0}|c)$ as:
$\pi^{*}(x_{0}|c)\propto\pi_{0}(x_{0}|c)\exp\left(\frac{1}{\beta}r(x_{0},c)\right),$ (4)
This aligns with Direct Preference Optimization (DPO) [27] that translates the optimization goal into training the optimal policy $\pi^{*}(x_{0}|c)$. The key advantage of this formulation is that it incorporates both reward maximization and regularization into a unified framework, thereby enabling robust learning. To improve the model’s alignment with human feedback and output quality, we incorporate real-world data into the alignment training process. By using authentic, human-generated examples, the model learns to reduce hallucinations and better align with human preferences. This also enhances the model’s robustness and generalization, preventing training divergence and ensuring reliable performance in diverse real-world tasks.
3.3.2 Pointwise Optimization
For pointwise preference alignment, we adopt Diffusion-KTO [7]. Unlike traditional preference-based methods, KTO optimizes a nonlinear utility function using binary-labeled data, making it particularly sensitive to negative samples. This approach avoids the need for explicit preference pairs and introduces human-aware loss functions.
In Diffusion-KTO, the utility of an action is computed relative to a reference point $Q_{\text{ref}}$, defined as the KL divergence between the learned policy $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$. The optimization objective is:
@@ 3.3.2 Pointwise Optimization
$L_{\text{kto}}(\theta)=\underset{\pi_{\theta}}{\text{max}}\mathbb{E}_{x_{0}\sim\mathcal{D},t\sim\text{Uniform([0,T])}}\\ [U(w(x_{0})(\beta\log\frac{\pi_{\theta}(x_{t-1}|x_{t})}{\pi_{\text{ref}}(x_{t-1}|x_{t})}-Q_{\text{ref}}))],$ (5)
where $w(x_{0})=\pm 1$ indicates whether a state $x_{0}$ is desirable or undesirable. The reference point $Q_{\text{ref}}$ is defined as:
$Q_{\text{ref}}=\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}(a|s)||\pi_{\text{ref}}(a|s)].$ (6)
This adaptive reference mechanism allows Diffusion-KTO to calibrate the policy toward human-aligned decision-making. To optimize the policy, we use the loss function from Eq. 5. A challenge in KTO training is estimating the average reward $Q$, which becomes unstable with small batch sizes. To mitigate this, we set the batch size as 128 for stable $Q$ estimation. To further improve stability and generalization, we incorporate real-world data from the VidGen-1M dataset [33]. We add a negative log-likelihood (NLL) loss term on real samples to regularize the model and prevent divergence from natural video distributions. The final training objective is:
@@
$\mathcal{L}(\theta)=L_{\text{kto}}(\theta)+\lambda\cdot\mathbb{E}_{(\textbf{{y}},x)\sim\mathcal{D}^{\mathrm{real}}}\bigl{[}-\log p_{\theta}(y|x)\bigr{]},$ (7)
where $\lambda$ controls the weight of real data regularization. This hybrid approach balances the model’s utility maximization with adherence to real-world video distributions, enhancing its alignment with human preferences.
3.3.3 Pairwise Optimization
For pairwise preference alignment, we adopt Diffusion-DPO [36], an extension of Direct Preference Optimization (DPO) [27] to diffusion models. Diffusion-DPO utilizes inverse decomposition to replace the standard policy $\pi_{\theta}$ and reference policy $\pi_{\text{ref}}$, enabling efficient preference optimization within the diffusion framework.
In Diffusion-DPO, the forward process $q(x_{1:T}|x_{0})$ approximates the reverse process $p_{\theta}(x_{1:T}|x_{0})$, redefining data likelihood in diffusion models. This leads to a loss function that facilitates robust preference optimization. The objective function for Diffusion-DPO is as:

### v4

# Dual-IPO: Dual-Iterative Preference Optimization for Text-to-Video Generation

@@ Abstract
Recent advances in video generation have enabled thrilling experiences in producing realistic videos driven by scalable diffusion transformers. However, they usually fail to produce satisfactory outputs that are aligned to users’ authentic demands and preferences. In this work, we introduce Dual-Iterative Optimization (Dual-IPO), an iterative paradigm that sequentially optimizes both the reward model and the video generation model for improved synthesis quality and human preference alignment. For the reward model, our framework ensures reliable and robust reward signals via CoT-guided reasoning, voting-based self-consistency, and preference certainty estimation. Given this, we optimize video foundation models with guidance of signals from reward model’s feedback, thus improving the synthesis quality in subject consistency, motion smoothness and aesthetic quality, etc. The reward model and video generation model complement each other and are progressively improved in the multi-round iteration, without requiring tediously manual preference annotations. Comprehensive experiments demonstrate that the proposed Dual-IPO can effectively and consistently improve the video generation quality of base model with various architectures and sizes, even help a model with only 2B parameters surpass a 5B one. Moreover, our analysis experiments and ablation studies identify the rational of our systematic design and the efficacy of each component.
Equal Contribution, *Corresponding Authors.Our source code is available at https://github.com/SAIS-FUXI/IPO.
@@ 1 Introduction
Video generation Brooks et al. (2024); Kuaishou (2024); hailuo (2024); Kong et al. (2024); Yang et al. (2024b); Tan et al. (2025) has experienced tremendous advancement in recent years. Their technical breakthroughs, driven by scalable model architectures (e.g., diffusion transformer Peebles and Xie (2023)), web-scale data and massive computation, have revolutionized traditional concent creation in various domains like movie production Zhang et al. (2025), story Zhou et al. (2024) and scene generation Guo et al. (2025), etc. Despite their success, existing video generation models still fall short in producing user-satisfied outputs with consistent subject, smooth motion and compelling aesthetic. This identifies a considerable gap between the quality of current models and the authentic demands of users in practice.
Figure 1: Overview of our proposed dual-iterative preference optimization framework. (a) Human preference dataset used to initially train our reward model. (b) Reward model that learns to automatically annotate generated videos with preference labels. (c) Dual-iterative optimization, which incorporates the reward model to iteratively optimize video foundation model, thus improving the synthesis quality and human alignment.
To improve synthesis quality and alignment of human preference, post-training techniques like Direct Preference Optimization (DPO) Rafailov et al. (2023); Liu et al. (2025b) and Kahneman-Tversky Optimization (KTO) Ethayarajh et al. (2024) are employed. However, preference learning typically requires training on large annotated datasets of human preferences, which are laborious to construct. Alternatively, external reward models, such as HPS Wu et al. (2023), ImageReward Xu et al. (2023), VideoAlign Liu et al. (2025a) and VisionReward Xu et al. (2024), can be utilized to evaluate the quality of images and videos for preference annotations. Nonetheless, these reward models often exhibit a distribution mismatch Kim et al. (2024) across different video generation models and evaluation criteria, making the reward signals unreliable for human preference alignment (see Tab. 1). Additionally, aligning video generation models on fixed offline preference datasets could easily cause overfitting and even collapsed to the preference signals Pal et al. (2024); Feng et al. (2024), degrading the synthesis quality and model generalization Park et al. (2024); Chen et al. (2024b); Azar et al. (2024). Furthermore, the reward signal should evolve with the model since the artifacts are becoming subtler throughout the training, otherwise the reward might be biased and lead to instable optimization Dubey et al. (2024).
To address the aforementioned challenges, this paper proposes Dual-iterative preference optimization (Dual-IPO), a post-training framework for improved synthesis quality and human alignment via iterative optimization on both the reward model and the video generation model. For the iterative refinement of our reward model, we propose self-refined preference optimization to unlock stronger reasoning capabilities from Vision-Language Models (VLMs) with a small set of Chain-of-Thought (CoT) guided annotated data, enabling high-quality and efficient reward modeling with minimal supervision. Then, we design a voting-based self-consistency mechanism across multiple inference paths to improve the accuracy and scalability of labeled rewards. Furthermore, a preference certainty estimator is developed to filter low-confidence results, ensuring reliable pseudo-labels for further training of the reward model. Such iterative self-labeling cycle strengthens the reward model’s generalization ability and robustness over time, ensuring reliable rewards for training video generation models. Regarding the iterative optimization of video generation models, the latest model is used to generate new videos in each iteration. These generated videos are scored by the current reward model, serving as updated reward signals for the next optimization iteration. Notably, our reward model supports both pairwise and pointwise preference annotation, facilitating flexible deployment under different optimization settings, i.e., pairwise for Diffusion-DPO and pointwise for Diffusion-KTO. The two component complement each other and together enhance the overall quality and user preference alignment of generated outputs. Besides, only a small amount of annotated preferences are required as cold start data for training our reward model, suggesting satisfactory data efficiency.
To testify the effectiveness of our proposed framework, we conduct extensive experiments on video generation models with different architectures (i.e., cross-attention DiT based Wanx Wan et al. (2025) and MMDiT based CogVideo Yang et al. (2024b)), different model scales (i.e., 2B and 5B). Experimental results demonstrate that our method reflects substantial performance gains across different baseline models in various aspects, including subject consistency, motion smoothness, visual aesthetic, etc. Moreover, the synthesis performance could be consistently improved with our proposed multi-round iterative optimization, indicating that both the reward model and the video generation model are progressively refined through each iteration. To sum up, our primary contributions are summarized blow:
  • A dual-iterative preference optimization framework for improving video generation models. By updating the reward model and the video generation model in a multi-round manner, we successfully improve the synthesis quality and human preference alignment.
  • A self-refined preference optimization strategy for reliable reward models. With CoT-guided reasoning, voting-based self-consistency, and preference certainty estimation, we ensure high-confidence and accurate reward signals iteratively, thus improving the generalization and robustness of the reward model. Notably, both pairwise and pointwise rewards are supported, enabling flexible deployment for different alignment strategies.
  • A multi-round scheme for consistently improvement on video quality. Leveraging these reliable rewards, we refine T2V models iteratively across multiple aspects (semantic, motion, aesthetic), consistently leading to better alignment and stronger generalization.
[…]
@@ 2 Related Work
Text-to-video generation. Recent advances in video generation have been largely driven by diffusion models Ho et al. (2020); Sohl-Dickstein et al. (2015); Song et al. (2020); Zhang et al. (2022), achieving remarkable improvements in diversity, fidelity, and quality Chen et al. (2023); Esser et al. (2023); Ho et al. (2022b); Liu et al. (2023a); Ruan et al. (2023); Wang et al. (2025); Xing et al. (2023). Scaling model size and training data further enhance performance Ho et al. (2022a); Singer et al. (2022); Wu et al. (2021). In text-to-video (T2V) generation, pretraining from scratch requires substantial computational resources and large-scale datasets Gen-3 (2024); Luma (2024); Kuaishou (2024); Brooks et al. (2024); Kong et al. (2024). Current methods Blattmann et al. (2023); Wang et al. (2023); Guo et al. (2023) extend text-to-image models by incorporating spatial and temporal attention modules, often using joint image-video training Ma et al. (2024), with Guo et al. (2023) proposing a plug-and-play temporal module for personalized image models. Remarkable visual quality and temporal consistency have been achieved by Zhang et al. (2023); Chen et al. (2024a), and diffusion transformers (DiT) Brooks et al. (2024); Yang et al. (2024b); Chen et al. (2024a) enhance spatiotemporal coherence and scalability, identifying the great potential of transformer architectures. Despite their success, they often fail to produce user-preferred videos in practice.
Post-training for human preference alignment. Post-training is widely adopted in Large Language Models (LLMs)Ouyang et al. (2022); Ramamurthy et al. (2022); Zheng et al. (2023). RLHFOuyang et al. (2022) aligns models via reward functions learned from comparisons. RRHF Yuan et al. (2023b) extends supervised fine-tuning, and SLiC Zhao et al. (2023) uses offline RL data. DPO Rafailov et al. (2023) improves stability by directly optimizing preference data, while KTO Ethayarajh et al. (2024) adopts a prospect-theoretic view. RSO Liu et al. (2023b) leverages rejection sampling for better alignment. Post-training has also been applied to image foundation models Clark et al. (2023); Prabhudesai et al. (2024a); Xu et al. (2023). For diffusion models, Diffusion-DPO Wallace et al. (2023) extends DPO to optimize likelihood, while Diffusion-KTO Li et al. (2025) maximizes expected utility without pairwise preferences. However, post-training for video generation remains underexplored. Existing efforts focus on reward design. InstructVideo Yuan et al. (2023a) proposes temporal decaying rewards (TAR) to prioritize central frames. T2V-Turbo Li et al. (2024) improves ODE solvers using motion priors. VADER Prabhudesai et al. (2024b) fine-tunes with vision-based reward gradients, boosting generalization. Despite these, gains remain limited, and the full potential of post-training is yet to be realized. In this paper, we propose a dual-optimization strategy to iteratively update the reward model and the video generation model. They complement each other and together deliver improved output quality and preference alignment.
@@ 3 METHOD
3.1 Method Overview
Fig. 1 presents the overall framework of our method. First, we collect a small amount of annotated user preferences for initially training our reward model. In this phase, both pairwise and pointwise annotations are constructed, enabling flexible alignment strategies like DPO with pairwise data and KTO with pointwise data. After training on the human preference dataset, we incorporate the reward model and the T2V generation model in the iterative optimization loop, performing dual-optimization to fully unlock their potentials. We detail each component of our method below.
3.2 Self-Refined Preference Optimization
We propose self-refined preference optimization (SRPO), a self-iterative training loop that progressively enhances the reward model with minimal supervision.
CoT-guided annotation.
We start by constructing a small set of Chain-of-Thought (CoT) guided annotations, unlocking the potential of a Vision-Language Model (VLM) for structured reasoning and reliable preference construction. Specifically, we employ CogVideoX-2B model Yang et al. (2024b) to generate videos from a diverse prompts built by Qwen2.5 Yang et al. (2024a). These prompts span over ten semantic categories (e.g., Human, Vehicle, Architecture) and multiple motion types (details are in the appendix). Each prompt yields four videos from different seeds, and the dataset is iteratively expanded using an optimized regeneration strategy. We collect both pairwise ranked and pointwise scored preference data along three dimensions: text-video consistency, content faithfulness, and motion smoothness following Tan et al. (2024b).
Self-consistency via multi-path inference.
For preference queries with deterministic answers, voting across multiple reasoning paths has been shown to improve reliability Prasad et al. (2024); Wang et al. (2022). Following this, SRPO performs multi-path inference with CoT-enabled reward model to scale the preference data. Concretely, we adopt a self-consistency mechanism that aggregates responses by counting answer frequencies, automatically constructing stable pseudo preference labels and reducing noise from individual paths.
Preference Certainty Estimator(PCE).
While self-consistency improves pseudo-label quality, noise can still arise due to randomness in VLM outputs and lack of reasoning trace inspection. To further filter unreliable labels, we introduce a preference certainty estimator (PCE) based on the reward formulation used in DPO Rafailov et al. (2023). Given a prompt $x$, let $y_{w}$ and $y_{l}$ denote the preferred and less-preferred videos in a pair. Formally, DPO models the preference probability Zhu et al. (2025); Kim et al. (2024) as:
@@ 3.2 Self-Refined Preference Optimization
$\mathbb{P}_{\theta}(y_{w}\!\succ\!y_{l}\!\mid\!x)\!=\!\!\sigma\!\left(\!\beta\!\!\log\!\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}\!-\!\!\beta\!\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\!\right),$ (1)
where $\pi_{\theta}$ and $\pi_{\text{ref}}$ are the trained and reference models, respectively, and $\beta$ is a temperature parameter. After DPO learning, the model implicitly learns a reward-like scoring to indicate the model’s preference $y$ over the baseline:
$R(y\mid x)=\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}.$ (2)
We define the average reward across all samples in the current dataset as $Q=\mathbb{E}_{(x,y)}[R(y\mid x)]$. To estimate the confidence that $y_{w}$ is truly better than the average, we define our certainty metric as:
$\text{PCE}(y_{w}\mid x)=\mathbb{P}_{\theta}\left(R(y_{w}\mid x)>Q\right).$ (3)
Intuitively, this measures how likely the chosen sample is better than a typical sample in the current distribution. We filter out weak or uncertain preferences by only retaining data points where $\text{PCE}(y_{w}\mid x)>0.5$. This criterion ensures that the selected preference aligns not just with relative ranking, but also exceeds the global quality threshold, improving label robustness under distribution shift.
Iterative Pseudo-Label Refinement.
The refined pseudo-labels are used to train the critic model, which is then applied to newly generated video samples for the next iteration. This bootstrapping process allows the critic to progressively improve its alignment with human preferences and generalize to evolving data distributions.
We adopt the DPO objective Rafailov et al. (2023), which increases the preference margin between the preferred sample $y_{w}$ and the less-preferred $y_{l}$ given the same input $x$ (e.g., text prompt). The training loss is defined as:
@@
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{x,y_{w},y_{l}}\bigg[$ $\log\sigma\bigg(\beta\log\frac{p_{\theta}(y_{w}\mid x)}{p_{\mathrm{ref}}(y_{w}\mid x)}-\beta\log\frac{p_{\theta}(y_{l}\mid x)}{p_{\mathrm{ref}}(y_{l}\mid x)}\bigg)\bigg]$
where $\sigma(\cdot)$ is the sigmoid function, $\beta$ is a temperature parameter, and $p_{\mathrm{ref}}$ is a frozen reference model. To enhance robustness, we further introduce the preference certainty estimator (PCE), defined as:
$\text{PCE}(y_{w}\mid x)=\mathbb{P}_{\theta}\left(R(y_{w}\mid x)>Q\right),$ (4)
where $R(y\mid x)=\log\frac{p_{\theta}(y\mid x)}{p_{\mathrm{ref}}(y\mid x)}$ is the learned reward score, and $Q=\mathbb{E}_{x,y}[R(y\mid x)]$ is the mean reward over the dataset. Consequently, the final loss incorporates PCE as a confidence weight:
@@
$\mathcal{L}_{\mathrm{SRPO}}(\theta)=\mathbb{E}_{x,y_{w},y_{l}}\left[\text{PCE}(y_{w}\mid x)\cdot\mathcal{L}_{\mathrm{DPO}}(\theta)\right].$ (5)
This formulation emphasizes reliable preferences during training and suppresses noisy supervision, leading to a more stable and generalizable reward model.
3.3 Iterative Alignment for Video Generation
After building an reliable reward model, we turn to align text-to-video (T2V) models with its feedback. Unlike one-time alignment methods that rely on offline fixed preference datasets, we iteratively leverage dynamic feedback from the self-improving reward model to provide robust and adaptive preference supervision throughout training.
Optimization loop.
Our core design is an iterative training loop between the T2V model and the reward model. In each iteration, the generator produces a batch of videos from diverse prompts. The reward model then evaluates these videos and produce pairwise data for DPO optimization and pointwise data for KTO optimization. To ensure training stability and efficient convergence, we monitor key indicators including the VBench metrics and training loss curves throughout the training process. When performance degradation (e.g., drop in semantic fidelity or motion quality) or abnormal loss dynamics are detected, the framework automatically triggers a critic update via the self-refined preference optimization procedure. This ensures more accurate and distribution-aware reward model, which in turn to re-score generated videos and reconstruct updated preference datasets for the next iteration. In this way, we can effectively mitigate reward misalignment and overfitting, enabling progressive preference-driven refinement of the T2V model.
Different preference strategies.
We support both pairwise and pointwise alignment paradigms. For pairwise supervision, we adopt Diffusion-DPO Wallace et al. (2023), which extends Direct Preference Optimization to the diffusion setting. Given preference pairs $(x_{0}^{w},x_{0}^{l})$ scored by the critic, Diffusion-DPO minimizes the following loss: